OpenStudy (anonymous):

factorise 10root2 x^2 + 3x - 9root2

OpenStudy (anonymous):

\[10\sqrt{2} x ^{2}+3x-9\sqrt{2}\]

OpenStudy (anonymous):

\[10\sqrt{2}x^2+3x-9\sqrt{2}\]?

OpenStudy (anonymous):

yes pls...

OpenStudy (anonymous):

ok i would not want to so this by had, so lets cheat

OpenStudy (anonymous):

how???

OpenStudy (anonymous):

set = 0 and solve using quadratic formula. get the zeros of \[r_1,r_2\] and then write \[(x-r_1)(x-r_2)\]

OpenStudy (anonymous):

the zeros are \[\frac{3\sqrt{2}}{5}\] and \[-\frac{3}{2\sqrt{2}}\]

OpenStudy (anonymous):

i get (-3+3root(89))/20root2 and (-3-3root(89))/20root2 as the roots... :/

OpenStudy (anonymous):

today i am too tired and brain-dead and did not feel like doing it as it baffled me a litttle so i put it out for u brainiacs to do it for me ... o( :{)

OpenStudy (anonymous):

wow...i'll have to check my math. mine are not pretty but satellite's are

OpenStudy (anonymous):

so we can factor as \[(5x-3\sqrt{2})(2\sqrt{2}x+3)\]

OpenStudy (anonymous):

i cheated. found the zeros, then factored. that is probably easiest way in this case. otherwise i am stuck

OpenStudy (anonymous):

you can check by multiplication that this works. think you can even do it in your head.

OpenStudy (anonymous):

first is right. last is right

OpenStudy (anonymous):

Guys, the student who brought this to me has to solve it by "splitting the middle term" method as he has not been taught any other method yet......

OpenStudy (anonymous):

well then lets split it.

OpenStudy (anonymous):

All DETAILED help appreciated today !!!!!

OpenStudy (anonymous):

$%^&* really?

OpenStudy (anonymous):

what is the gimmick? we want to split the middle term. guess maybe we can work backwards and see what it should be

OpenStudy (anonymous):

see what i meant... its a foxy one !!!

myininaya (myininaya):

3x=15x-12x try that

OpenStudy (anonymous):

my hero!

OpenStudy (anonymous):

still not clear how it works though. can you elaborate?

OpenStudy (anonymous):

yeah that is the crux i cud not work out finding factors of 180 was no problem further steps eluded me

OpenStudy (anonymous):

now i am intrigued. can you explain?

OpenStudy (anonymous):

although i have to admit i like my way better!

OpenStudy (anonymous):

i think i got it......

myininaya (myininaya):

(sqrt(2))^4=4

OpenStudy (anonymous):

\[10\sqrt{2}x ^{2} + 15x - 12x -9\sqrt{2}\]

myininaya (myininaya):

\[10\sqrt{2}x^2+15x-12x-9\sqrt{2}=5x(2\sqrt{2}x+3)-3(4x+3\sqrt{2})=5x(2\sqrt{2}x+3)-3((\sqrt{2})^4x+3\]

OpenStudy (anonymous):

\[5x(2\sqrt{2}x +3) -3\sqrt{2}(2\sqrt{2}x + 3)\]

myininaya (myininaya):

\[5x(2\sqrt{2}x+3)-3((\sqrt{2})^4x+3\sqrt{2})=5x(2\sqrt{2}x+3)-3\sqrt{2}(2\sqrt{2}x+3)\]

myininaya (myininaya):

yes! and the rest is easy

OpenStudy (anonymous):

\[(5x-3\sqrt{2})(2\sqrt{2}x + 3)\]

OpenStudy (anonymous):

@Harkiat. I was almost there...i went the other way and pulled the 2root(2)x + 3 out from each :)

OpenStudy (anonymous):

pls check my working if its okay???

myininaya (myininaya):

thats what i got above

myininaya (myininaya):

gj

OpenStudy (anonymous):

I get the same answer...not sure i'd be able to motivate the split without some trouble though.

OpenStudy (anonymous):

@myininaya thanks for the good work !! ALL OTHERS - Thanks for getting my juices flowing again. Combined "brain-storming" defeats the PROBLEM once more !!!!!

OpenStudy (anonymous):

lorda mercy i am sticking to finding the zeros, or at the very least completing the square. god work though!

OpenStudy (anonymous):

i mean good work. maybe god's work too, who knows?

myininaya (myininaya):

i like this is a good bonus question for peeps in college algebra

OpenStudy (anonymous):

sure if you want them to really hate you!

myininaya (myininaya):

lol they wouldn't get it

OpenStudy (anonymous):

JAI HO !!! ( as in the song from SLUMDOG MILLIONAIRE)

OpenStudy (anonymous):

MEDALS to ALL - every thread of contribution appreciated, with special word of praise for myininaya (Is that yr real name???) !!!!!!

myininaya (myininaya):

no it isn't i want to call my daughter myininaya

myininaya (myininaya):

i made it up im in love with the name

OpenStudy (anonymous):

what does it mean???

myininaya (myininaya):

i guess i could make up a meaning lol

OpenStudy (anonymous):

where r u from, I am from Delhi - India

OpenStudy (anonymous):

@ satellite CONGRATS FOR CROSSING 1000 mark !!!!

myininaya (myininaya):

united states